OpenStudy (anonymous):

MEDAL & FAN How far from the origin is the point (-5, -12)?

OpenStudy (jadeishere):

Do you kow what the Pythagoras Theorem is?

OpenStudy (anonymous):

my parents are making me take high school classes over the summer and I'm in 6th grade, so to answer your question I do not.

OpenStudy (jadeishere):

Okay :P Well, the Pythagoras Theorem is a formula a^2 + b^2 = c^2 a and b are the two legs, and in this case, a would be -5 and b = -12 (or vice versa) The distance would be C, that's called the hypotenuse of a right triangle. so -5 squared is 5, since multiplying a negative by another negative equals a positive. and -12 squared is 144 so 144 + 25 = 169 then you find the square root of 169 and you have your answer. the square root of 169 is 13 Hope this helps :)

OpenStudy (jadeishere):

Well, okey. Have a good day :) And welcome to Openstudy :P